OpenStudy (anonymous):

why is this the complete ground state electron configuration of a phosphorous atom? 1s^2 2s^2 2p^3 3s^2 3p^3 I was taught the numbers in the exponents must equal the number of electrons and these numbers do not add up

OpenStudy (rogue):

There's a mistake with the 2p orbitals for the config. you have. Phosphorus has 15 electrons in the ground state, so its configuration should be\[P: 1s^2 2s^2 2p^6 3s^2 3p^3\]

OpenStudy (anonymous):

but the powers add up to 13

OpenStudy (anonymous):

shouldnt they add up to 15?

OpenStudy (rogue):

Yes, they should. The one that the textbook/techer gave you is wrong.

OpenStudy (anonymous):

thanks for letting me know. can you work with me please? Im preparing for a test. I have some more questions

OpenStudy (anonymous):

calculate the frequency of red light at 650nm

OpenStudy (anonymous):

the correct answer to this is 4.6 x 10^14 s^-1 but i dont see how. I got different exponents

OpenStudy (rogue):

That's an easy one.\[c = f \lambda\]c is a constant, its the speed of light, which is approx 3 x 10^8 m/s. f is frequency and lambda is wavelength.

OpenStudy (anonymous):

c/f right?

OpenStudy (anonymous):

but convert nm to m first

OpenStudy (rogue):

\[f = \frac {c}{\lambda}\]And yes, convert to meters first.

OpenStudy (anonymous):

3.00 x 10^8/ 6.5 x 10^-7

OpenStudy (anonymous):

i just get 4.61

OpenStudy (anonymous):

no exp

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

actually explain sorry

OpenStudy (rogue):

:P negative exp

OpenStudy (anonymous):

huh

OpenStudy (rogue):

I think you entered it in the calculator wrong

OpenStudy (anonymous):

yea i figured thanks

OpenStudy (rogue):

Alright, np =)

OpenStudy (anonymous):

while we are on frequency... A photon has a frequency of 6.5 x 10^3 MHz. Convert this frequency into wavelength, in nm

OpenStudy (rogue):

Same equation, convert MHz to Hz; solve for lambda. Hertz is just s^-1 btw.

OpenStudy (anonymous):

6500000 Hz

OpenStudy (anonymous):

3.00 x 10^8/6500000 = 46.153m

OpenStudy (anonymous):

46.153m x 10^9nm= 4.615 x 10^10....the correct exp is suppose to be 7

OpenStudy (rogue):

The freq should be 6.5 x 10^9 Hz

OpenStudy (anonymous):

6.5 x 10^3 (1000)?

OpenStudy (anonymous):

thats what i entered and it gave me 6500000

OpenStudy (rogue):

6.5 x 10^3 MHz = 6.5 x 10^3 x 10^6 Hz = 6.5 GHz = 6.5 x 10^9 Hz

OpenStudy (rogue):

6500 megahertz to hertz...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

whats after that

OpenStudy (anonymous):

c/f?

OpenStudy (rogue):

yeah.

OpenStudy (anonymous):

3.00 x 10^8/ 6.5 x 10^9

OpenStudy (anonymous):

??

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

another question

OpenStudy (anonymous):

Place the following in order of decreasing radius Br^(-), Kr, Se^(2-), Sr^(2+), Rb^(+)

OpenStudy (rogue):

Br^- has 35 protons and 36 electrons. Kr has 36 protons and 36 electrons. Rb^+ has 37 protons and 36 electrons. Sr^(2+) has 38 protons and 36 electrons. Think about the effective nuclear charge. The stronger the pull of the nucleus, the closer in electrons will go. All of the have the same # of electrons, but different # of protons. The one with the highest # of protons has the strongest pull and so the smallest radius. So the order in decreasing radius is Br^-, Kr, Rb^+, Sr^(2+)

OpenStudy (anonymous):

wow thanks for that brilliant explanation

OpenStudy (anonymous):

the electron spin quantum number has two possible values +1/2 and -1/2...how do you determine the value for instance when n=3 l=0 Ml=0 Ms= +1/2 and when n=3 l=0 Ml=0 Ms= -1/2